# Finite-Time Analysis of Simultaneous Double Q-learning

### v1

@@ B.4 CONVERGENCE OF $Q_{k}^{\text{err}}$
$\mathbb{E}\left[\lVert Q_{k}^{\text{err}}\rVert_{\infty}\right]$ $\leq\frac{8\gamma d_{\text{max}}\lvert S\times A\rvert\alpha^{1/2}}{d_{\text{min}}^{5/2}(1-\gamma)^{7/2}}+\frac{8\alpha^{1/2}\lvert S\times A\rvert}{d_{\text{min}}^{3/2}(1-\gamma)^{5/2}}+\frac{4\rho^{k-2}k^{2}\alpha\gamma d_{\text{max}}\lvert S\times A\rvert^{3/2}}{(1-\gamma)}$
$\quad+\frac{4k\rho^{k-1}\lvert S\times A\rvert^{3/2}}{(1-\gamma)}.$ (28)
@@
$\leq\rho^{k}\frac{4\gamma^{2}d_{\text{max}}^{2}\lvert S\times A\rvert\alpha^{1/2}}{d_{\text{min}}^{5/2}(1-\gamma)^{7/2}}+\rho^{k}\frac{4\gamma d_{\text{max}}\lvert S\times A\rvert\alpha^{1/2}}{d_{\text{min}}^{3/2}(1-\gamma)^{5/2}}$
$\quad+\rho^{k-2}(k-1)(k-2)\frac{\alpha^{2}\gamma^{2}d_{\text{max}}^{2}\lvert S\times A\rvert^{3/2}}{1-\gamma}$
$\quad+k\rho^{k-1}2\alpha\gamma d_{\text{max}}\frac{\lvert S\times A\rvert^{3/2}}{1-\gamma}+\frac{4\gamma d_{\text{max}}\lvert S\times A\rvert\alpha^{1/2}}{d_{\text{min}}^{3/2}(1-\gamma)^{5/2}}$
$\quad+k\rho^{k-1}2\alpha\gamma d_{\text{max}}\frac{\lvert S\times A\rvert^{3/2}}{1-\gamma}$
$\quad+\frac{4\alpha^{1/2}\lvert S\times A\rvert}{d_{\text{min}}^{1/2}(1-\gamma)^{3/2}}+\lvert S\times A\rvert^{3/2}\frac{2}{1-\gamma}\rho^{k}$ (29)
Then we group some terms of (B.4) as
$\rho^{k}\frac{4\gamma^{2}d_{\text{max}}^{2}\lvert S\times A\rvert\alpha^{1/2}}{d_{\text{min}}^{5/2}(1-\gamma)^{7/2}}+\rho^{k}\frac{4\gamma d_{\text{max}}\lvert S\times A\rvert\alpha^{1/2}}{d_{\text{min}}^{3/2}(1-\gamma)^{5/2}}\leq\frac{8\gamma d_{\text{max}}\lvert S\times A\rvert\alpha^{1/2}}{d_{\text{min}}^{5/2}(1-\gamma)^{7/2}}$
$\frac{4\gamma d_{\text{max}}\lvert S\times A\rvert\alpha^{1/2}}{d_{\text{min}}^{3/2}(1-\gamma)^{5/2}}+\frac{4\alpha^{1/2}\lvert S\times A\rvert}{d_{\text{min}}^{1/2}(1-\gamma)^{3/2}}\leq\frac{8\alpha^{1/2}\lvert S\times A\rvert}{d_{\text{min}}^{3/2}(1-\gamma)^{5/2}},$
$\rho^{k-2}(k-1)(k-2)\frac{\alpha^{2}\gamma^{2}d_{\text{max}}^{2}\lvert S\times A\rvert^{3/2}}{1-\gamma}+k\rho^{k-1}2\alpha\gamma d_{\text{max}}\frac{\lvert S\times A\rvert^{3/2}}{1-\gamma}$
$\leq\frac{4\rho^{k-2}k^{2}\alpha\gamma d_{\text{max}}\lvert S\times A\rvert^{3/2}}{(1-\gamma)},$
$\lvert S\times A\rvert^{3/2}\frac{2}{1-\gamma}\rho^{k}+k\rho^{k-1}2\alpha\gamma d_{\text{max}}\frac{\lvert S\times A\rvert^{3/2}}{1-\gamma}\leq\frac{4k\rho^{k-1}\lvert S\times A\rvert^{3/2}}{(1-\gamma)}$
@@
$\mathbb{E}\left[\lVert Q_{k}^{\text{err}}\rVert_{\infty}\right]$ $\leq\frac{8\gamma d_{\text{max}}\lvert S\times A\rvert\alpha^{1/2}}{d_{\text{min}}^{5/2}(1-\gamma)^{7/2}}+\frac{8\alpha^{1/2}\lvert S\times A\rvert}{d_{\text{min}}^{3/2}(1-\gamma)^{5/2}}+\frac{4\rho^{k-2}k^{2}\alpha\gamma d_{\text{max}}\lvert S\times A\rvert^{3/2}}{(1-\gamma)}$
$\quad+\frac{4k\rho^{k-1}\lvert S\times A\rvert^{3/2}}{(1-\gamma)}$
@@ Appendix C Detailed analysis result of convergence of SDQ (Remaining part)
In this section, convergence of SDQ will be studied based on the results in Section B. In Section B, a bound on $Q_{k}^{\text{err}}$ has been obtained in Lemma 4. To compete the proof, the following three steps remain:
@@
where the first inequality is due to $\Pi_{Q^{*}}Q^{*}\geq\Pi_{Q_{k}^{B}}Q^{*}$, and the second inequality is due to the hypothesis $Q_{k}^{A_{U}}-Q^{*}\geq Q_{k}^{A}-Q^{*}$. And the second inequality is due to the hypothesis $Q_{k}^{A_{U}}\geq Q_{k}^{A}$ and the fact that the matrix $I+\alpha\gamma DP\Pi_{Q_{k}^{B}}-\alpha D$ is nonnegative, i.e., all elements are nonnegative by Lemma 1. Therefore, by induction argument, one concludes $Q_{k}^{A_{U}}-Q^{*}\geq Q_{k}^{A}-Q^{*}$ for all $k\geq 0$. The proof of the second inequality follows similar lines. This completes the proof.
@@
$\begin{split}Q_{k+1}^{A}-Q^{*}&=Q_{k}^{A}-Q^{*}+\alpha D\{\gamma P\Pi_{Q_{k}^{B}}Q_{k}^{A}-\gamma P\Pi_{Q^{*}}Q^{*}-Q_{k}^{A}+Q^{*}\}+\alpha w_{k}^{A}\\ &=(I-\alpha D)(Q_{k}^{A}-Q^{*})+\alpha D\{\gamma P\Pi_{Q_{k}^{B}}Q_{k}^{B}-\gamma P\Pi_{Q^{*}}Q^{*}+\gamma P\Pi_{Q_{k}^{B}}(Q_{k}^{A}-Q_{k}^{B})\}\\ &\quad+\alpha w_{k}^{A}\\ &\geq(I-\alpha D)(Q_{k}^{A}-Q^{*})+\alpha D\{\gamma P\Pi_{Q^{*}}Q_{k}^{B}-\gamma P\Pi_{Q^{*}}Q^{*}+\gamma P\Pi_{Q_{k}^{B}}(Q_{k}^{A}-Q_{k}^{B})\}\\ &\quad+\alpha w_{k}^{A}\\ &=(I+\alpha\gamma DP\Pi_{Q^{*}}-\alpha D)(Q_{k}^{A}-Q^{*})+\alpha\gamma DP(\Pi_{Q_{k}^{B}}-\Pi_{Q^{*}})(Q_{k}^{A}-Q_{k}^{B})\\ &\quad+\alpha w_{k}^{A}\\ &\geq(I+\alpha\gamma DP\Pi_{Q^{*}}-\alpha D)(Q_{k}^{A_{L}}-Q^{*})+\alpha\gamma DP(\Pi_{Q_{k}^{B}}-\Pi_{Q^{*}})(Q_{k}^{A}-Q_{k}^{B})\\ &\quad+\alpha w_{k}^{A}\\ &=Q_{k+1}^{A_{L}}-Q^{*},\end{split}$
where the first inequality is due to $\Pi_{Q^{B}}Q^{B}\geq\Pi_{Q^{*}}Q^{B}$, and the second inequality is due to the hypothesis $Q_{k}^{A_{L}}-Q^{*}\leq Q_{k}^{A}-Q^{*}$. Therefore, by induction argument, one concludes $Q_{k}^{A_{L}}-Q^{*}\leq Q_{k}^{A}-Q^{*}$ for all $k\geq 0$. And the second inequality is due to the hypothesis $Q_{k}^{{A_{L}}}\leq Q_{k}^{A}$ and the fact that the matrix $I+\alpha\gamma DP\Pi_{Q{*}}-\alpha D$ is nonnegative, i.e., all elements are nonnegative by Lemma 1 The proof of the second inequality follows lines similar to the first proof. This completes the proof.
@@ C.3 Convergence of the lower comparison system
The lower comparison system in (5.3) can be divided into the linear parts with stochastic noises, $(I+\alpha\gamma DP\Pi_{Q^{*}}-\alpha D)(Q_{k}^{A_{L}}-Q^{*})+\alpha w_{k}^{A}$ and $(I+\alpha\gamma DP\Pi_{Q^{*}}-\alpha D)(Q_{k}^{B_{L}}-Q^{*})+\alpha w_{k}^{B}$, and the external disturbance parts, $\alpha\gamma DP(\Pi_{Q_{k}^{B}}-\Pi_{Q^{*}})(Q_{k}^{A}-Q_{k}^{B})$ and $\alpha\gamma DP(\Pi_{Q_{k}^{*}}-\Pi_{Q_{k}^{A}})(Q_{k}^{A}-Q_{k}^{B})$. As proved in Section B.4, the external disturbances are bounded. Using this fact, one can prove the finite-time error bounds of the linear part with stochastic noise as $(I+\alpha\gamma DP\Pi_{Q^{*}}-\alpha D)(Q_{k}^{A_{L}}-Q^{*})+\alpha w_{k}^{A}$ and $(I+\alpha\gamma DP\Pi_{Q^{*}}-\alpha D)(Q_{k}^{B_{L}}-Q^{*})+\alpha w_{k}^{B}$.
Theorem C.1.
@@
$\mathbb{E}[\lVert Q_{k}^{A_{L}}-Q^{*}\rVert_{\infty}]$ $\leq\frac{16\gamma d_{\text{max}}\lvert S\times A\rvert\alpha^{1/2}}{d_{\text{min}}^{7/2}(1-\gamma)^{9/2}}+\frac{24\rho^{k-3}k^{3}\lvert S\times A\rvert^{3/2}}{(1-\gamma)}+\frac{4\alpha^{1/2}\lvert S\times A\rvert}{d_{\text{min}}^{1/2}(1-\gamma)^{3/2}}.$ (30)
@@
$Q_{k}^{{A_{L}}}-{Q^{*}}=$ ${(I+\alpha\gamma DP{\Pi_{{Q^{*}}}}-\alpha D)^{k}}(Q_{0}^{{A_{L}}}-{Q^{*}})+\underbrace{\sum\limits_{j=0}^{k-1}\alpha{{(I+\alpha\gamma DP{\Pi_{{Q^{*}}}}-\alpha D)}^{(k-1)-j}}w_{j}^{A}}_{=:(*)}$
$+\underbrace{\alpha\gamma DP\sum\limits_{j=0}^{k-1}{{{(I+\alpha\gamma DP{\Pi_{{Q^{*}}}}-\alpha D)}^{(k-1)-j}}}({\Pi_{Q_{j}^{B}}}-{\Pi_{{Q^{*}}}})(Q_{j}^{A}-Q_{j}^{B})}_{=:(**)},$
where ()(*)( ∗ ) reflects the effect of the stochastic noise $w_{j}^{A}$ and ()(**)( ∗ ∗ ) corresponds to the effect of the disturbance $Q_{j}^{A}-Q_{j}^{B}$. Taking the $\infty$-norm on the right-hand side of the above equation leads to
$\lVert Q_{k}^{A_{L}}-Q^{*}\rVert_{\infty}=$ $\biggl{\lVert}(I+\alpha\gamma DP\Pi_{Q^{*}}-\alpha D)^{k}(Q_{0}^{A_{L}}-Q^{*})$
$\quad+\sum_{j=0}^{k-1}\alpha(I+\alpha\gamma DP\Pi_{Q^{*}}-\alpha D)^{(k-1)-j}w_{j}^{A}$
$+\alpha\gamma DP\sum_{j=0}^{k-1}(I+\alpha\gamma DP\Pi_{Q^{*}}-\alpha D)^{(k-1)-j}(\Pi_{Q_{j}^{B}}-\Pi_{Q^{*}})(Q_{j}^{A}-Q_{j}^{B})\biggr{\rVert}_{\infty}$
$\leq$ $\biggl{\lVert}(I+\alpha\gamma DP\Pi_{Q^{*}}-\alpha D)^{k}(Q_{0}^{A_{L}}-Q^{*})$
$\quad+\sum_{j=0}^{k-1}\alpha(I+\alpha\gamma DP\Pi_{Q^{*}}-\alpha D)^{(k-1)-j}w_{j}^{A}\biggr{\rVert}_{\infty}$
$+\biggl{\lVert}\alpha\gamma DP\sum_{j=0}^{k-1}(I+\alpha\gamma DP\Pi_{Q^{*}}-\alpha D)^{(k-1)-j}(\Pi_{Q_{j}^{B}}-\Pi_{Q^{*}})(Q_{j}^{A}-Q_{j}^{B})\biggr{\rVert}_{\infty}$
$\leq$ $\underbrace{\biggl{\lVert}(I+\alpha\gamma DP{\Pi_{{Q^{*}}}}-\alpha D)^{k}(Q_{0}^{{A_{L}}}-{Q^{*}})}_{:=(*)}$
$\underbrace{+\sum\limits_{j=0}^{k-1}\alpha{{(I+\alpha\gamma DP{\Pi_{{Q^{*}}}}-\alpha D)}^{(k-1)-j}}w_{j}^{A}\biggr{\rVert}_{\infty}}_{:=(**)}$
@@
where ()(*)( ∗ ) and ()(**)( ∗ ∗ ) in the second inequality corresponds to the solution of (15) with $x_{k}=Q_{k}^{{A_{L}}}-{Q^{*}}$ and $w_{k}=w_{k}^{A}$, we can apply the bound given in  Corollary A.3. Moreover, applying Lemma 3, one gets
$\lVert Q_{k}^{A_{L}}-Q^{*}\rVert_{\infty}\leq$ $\frac{4\alpha^{1/2}\lvert S\times A\rvert}{d_{\text{min}}^{1/2}(1-\gamma)^{3/2}}+\lvert S\times A\rvert^{3/2}\frac{2}{1-\gamma}\rho^{k}$
$\quad+\alpha\gamma d_{\text{max}}\sum_{j=0}^{k-1}\rho^{(k-1)-j}\biggl{\lVert}(\Pi_{Q_{j}^{B}}-\Pi_{Q^{*}})\biggr{\rVert}_{\infty}\biggl{\lVert}(Q_{j}^{A}-Q_{j}^{B})\biggr{\rVert}_{\infty},$
@@
$\mathbb{E}[\lVert Q_{k}^{A_{L}}-Q^{*}\rVert_{\infty}]$ $\leq\frac{8\gamma^{2}d_{\text{max}}^{2}\lvert S\times A\rvert\alpha^{1/2}}{d_{\text{min}}^{7/2}(1-\gamma)^{9/2}}+\frac{8\gamma d_{\text{max}}\lvert S\times A\rvert\alpha^{1/2}}{d_{\text{min}}^{5/2}(1-\gamma)^{7/2}}$
$\quad+\rho^{k-1}\biggl{(}\frac{(k-1)k(2k-1)}{6}\biggl{)}\frac{4\rho^{-2}\alpha^{2}\gamma^{2}d_{\text{max}}^{2}\lvert S\times A\rvert^{3/2}}{(1-\gamma)}$
$\quad+\rho^{k-1}\biggl{(}\frac{(k-1)k}{2}\biggl{)}\frac{4\rho^{-1}\alpha\gamma d_{\text{max}}\lvert S\times A\rvert^{3/2}}{(1-\gamma)}+\frac{4\alpha^{1/2}\lvert S\times A\rvert}{d_{\text{min}}^{1/2}(1-\gamma)^{3/2}}$
$\quad+\frac{2\rho^{k}\lvert S\times A\rvert^{3/2}}{(1-\gamma)}$ (31)
@@
$\frac{8\gamma^{2}d_{\text{max}}^{2}\lvert S\times A\rvert\alpha^{1/2}}{d_{\text{min}}^{7/2}(1-\gamma)^{9/2}}+\frac{8\gamma d_{\text{max}}\lvert S\times A\rvert\alpha^{1/2}}{d_{\text{min}}^{5/2}(1-\gamma)^{7/2}}\leq 2\biggl{(}\frac{8\gamma d_{\text{max}}\lvert S\times A\rvert\alpha^{1/2}}{d_{\text{min}}^{7/2}(1-\gamma)^{9/2}}\biggl{)}$

### v2

@@ B.4 CONVERGENCE OF $Q_{k}^{\mathrm{err}}$
$\mathbb{E}\left[\lVert Q_{k}^{\mathrm{err}}\rVert_{\infty}\right]\leq\frac{8\gamma d_{\mathrm{max}}\lvert\mathcal{S}\times\mathcal{A}\rvert\alpha^{1/2}}{d_{\mathrm{min}}^{5/2}(1-\gamma)^{7/2}}+\frac{8\alpha^{1/2}\lvert\mathcal{S}\times\mathcal{A}\rvert}{d_{\mathrm{min}}^{3/2}(1-\gamma)^{5/2}}+\frac{4\rho^{k-2}k^{2}\alpha\gamma d_{\mathrm{max}}\lvert\mathcal{S}\times\mathcal{A}\rvert^{3/2}}{(1-\gamma)}+\frac{4k\rho^{k-1}\lvert\mathcal{S}\times\mathcal{A}\rvert^{3/2}}{(1-\gamma)}.$ (29)
@@
$\leq\rho^{k}\frac{4\gamma^{2}d_{\text{max}}^{2}\lvert\mathcal{S}\times\mathcal{A}\rvert\alpha^{1/2}}{d_{\text{min}}^{5/2}(1-\gamma)^{7/2}}+\rho^{k}\frac{4\gamma d_{\text{max}}\lvert\mathcal{S}\times\mathcal{A}\rvert\alpha^{1/2}}{d_{\text{min}}^{3/2}(1-\gamma)^{5/2}}+\rho^{k-2}(k-1)(k-2)\frac{\alpha^{2}\gamma^{2}d_{\text{max}}^{2}\lvert\mathcal{S}\times\mathcal{A}\rvert^{3/2}}{1-\gamma}$
$\quad+k\rho^{k-1}2\alpha\gamma d_{\text{max}}\frac{\lvert\mathcal{S}\times\mathcal{A}\rvert^{3/2}}{1-\gamma}+\frac{4\gamma d_{\text{max}}\lvert\mathcal{S}\times\mathcal{A}\rvert\alpha^{1/2}}{d_{\text{min}}^{3/2}(1-\gamma)^{5/2}}+k\rho^{k-1}2\alpha\gamma d_{\text{max}}\frac{\lvert\mathcal{S}\times\mathcal{A}\rvert^{3/2}}{1-\gamma}$
$\quad+\frac{4\alpha^{1/2}\lvert\mathcal{S}\times\mathcal{A}\rvert}{d_{\text{min}}^{1/2}(1-\gamma)^{3/2}}+\lvert\mathcal{S}\times\mathcal{A}\rvert^{3/2}\frac{2}{1-\gamma}\rho^{k}$ (30)
$\rho^{k}\frac{4\gamma^{2}d_{\text{max}}^{2}\lvert\mathcal{S}\times\mathcal{A}\rvert\alpha^{1/2}}{d_{\text{min}}^{5/2}(1-\gamma)^{7/2}}+\rho^{k}\frac{4\gamma d_{\text{max}}\lvert\mathcal{S}\times\mathcal{A}\rvert\alpha^{1/2}}{d_{\text{min}}^{3/2}(1-\gamma)^{5/2}}\leq\frac{8\gamma d_{\text{max}}\lvert\mathcal{S}\times\mathcal{A}\rvert\alpha^{1/2}}{d_{\text{min}}^{5/2}(1-\gamma)^{7/2}},$
$\frac{4\gamma d_{\text{max}}\lvert\mathcal{S}\times\mathcal{A}\rvert\alpha^{1/2}}{d_{\text{min}}^{3/2}(1-\gamma)^{5/2}}+\frac{4\alpha^{1/2}\lvert\mathcal{S}\times\mathcal{A}\rvert}{d_{\text{min}}^{1/2}(1-\gamma)^{3/2}}\leq\frac{8\alpha^{1/2}\lvert\mathcal{S}\times\mathcal{A}\rvert}{d_{\text{min}}^{3/2}(1-\gamma)^{5/2}},$
$\rho^{k-2}(k-1)(k-2)\frac{\alpha^{2}\gamma^{2}d_{\text{max}}^{2}\lvert\mathcal{S}\times\mathcal{A}\rvert^{3/2}}{1-\gamma}+2k\rho^{k-1}\alpha\gamma d_{\text{max}}\frac{\lvert\mathcal{S}\times\mathcal{A}\rvert^{3/2}}{1-\gamma}\leq\frac{4\rho^{k-2}k^{2}\alpha\gamma d_{\text{max}}\lvert\mathcal{S}\times\mathcal{A}\rvert^{3/2}}{1-\gamma},$
$\frac{2\rho^{k}\lvert\mathcal{S}\times\mathcal{A}\rvert^{3/2}}{1-\gamma}+2k\rho^{k-1}\alpha\gamma d_{\text{max}}\frac{\lvert\mathcal{S}\times\mathcal{A}\rvert^{3/2}}{1-\gamma}\leq\frac{4k\rho^{k-1}\lvert\mathcal{S}\times\mathcal{A}\rvert^{3/2}}{1-\gamma}.$
@@
$\mathbb{E}\left[\lVert Q_{k}^{\text{err}}\rVert_{\infty}\right]$ $\leq\frac{8\gamma d_{\text{max}}\lvert\mathcal{S}\times\mathcal{A}\rvert\alpha^{1/2}}{d_{\text{min}}^{5/2}(1-\gamma)^{7/2}}+\frac{8\alpha^{1/2}\lvert\mathcal{S}\times\mathcal{A}\rvert}{d_{\text{min}}^{3/2}(1-\gamma)^{5/2}}+\frac{4\rho^{k-2}k^{2}\alpha\gamma d_{\text{max}}\lvert\mathcal{S}\times\mathcal{A}\rvert^{3/2}}{(1-\gamma)}+\frac{4k\rho^{k-1}\lvert\mathcal{S}\times\mathcal{A}\rvert^{3/2}}{(1-\gamma)}.$
This completes the finite-time error bound for $Q_{k}^{\text{err}}$ by combining the upper, lower, and upper–lower comparison systems.
@@ Appendix C Detailed analysis result of convergence of SDQ (Remaining part)
In this section, convergence of SDQ will be studied based on the results in B. In B, a bound on $Q_{k}^{\text{err}}$ has been obtained in Lemma 8. To compete the proof, the following three steps remain:
  • 1.
@@
where the first inequality is due to $\Pi_{Q^{*}}Q^{*}\geq\Pi_{Q_{k}^{B}}Q^{*}$, and the second inequality is due to the hypothesis $Q_{k}^{A_{U}}-Q^{*}\geq Q_{k}^{A}-Q^{*}$ and the fact that the matrix $I+\alpha\gamma DP\Pi_{Q_{k}^{B}}-\alpha D$ is nonnegative, i.e., all elements are nonnegative by Lemma 4. Therefore, by induction argument, one concludes $Q_{k}^{A_{U}}-Q^{*}\geq Q_{k}^{A}-Q^{*}$ for all $k\geq 0$. The proof of the second inequality follows similar lines. This completes the proof.
@@
$\begin{split}Q_{k+1}^{A}-Q^{*}&=Q_{k}^{A}-Q^{*}+\alpha D\{\gamma P\Pi_{Q_{k}^{B}}Q_{k}^{A}-\gamma P\Pi_{Q^{*}}Q^{*}-Q_{k}^{A}+Q^{*}\}+\alpha w_{k}^{A}\\ &=(I-\alpha D)(Q_{k}^{A}-Q^{*})+\alpha D\{\gamma P\Pi_{Q_{k}^{B}}Q_{k}^{B}-\gamma P\Pi_{Q^{*}}Q^{*}+\gamma P\Pi_{Q_{k}^{B}}(Q_{k}^{A}-Q_{k}^{B})\}+\alpha w_{k}^{A}\\ &\geq(I-\alpha D)(Q_{k}^{A}-Q^{*})+\alpha D\{\gamma P\Pi_{Q^{*}}Q_{k}^{B}-\gamma P\Pi_{Q^{*}}Q^{*}+\gamma P\Pi_{Q_{k}^{B}}(Q_{k}^{A}-Q_{k}^{B})\}+\alpha w_{k}^{A}\\ &=(I+\alpha\gamma DP\Pi_{Q^{*}}-\alpha D)(Q_{k}^{A}-Q^{*})+\alpha\gamma DP(\Pi_{Q_{k}^{B}}-\Pi_{Q^{*}})(Q_{k}^{A}-Q_{k}^{B})+\alpha w_{k}^{A}\\ &\geq(I+\alpha\gamma DP\Pi_{Q^{*}}-\alpha D)(Q_{k}^{A_{L}}-Q^{*})+\alpha\gamma DP(\Pi_{Q_{k}^{B}}-\Pi_{Q^{*}})(Q_{k}^{A}-Q_{k}^{B})+\alpha w_{k}^{A}\\ &=Q_{k+1}^{A_{L}}-Q^{*},\end{split}$
where the first inequality is due to $\Pi_{Q^{B}}Q^{B}\geq\Pi_{Q^{*}}Q^{B}$, and the second inequality is due to the hypothesis $Q_{k}^{A_{L}}-Q^{*}\leq Q_{k}^{A}-Q^{*}$. Therefore, by induction argument, one concludes $Q_{k}^{A_{L}}-Q^{*}\leq Q_{k}^{A}-Q^{*}$ for all $k\geq 0$. And the second inequality is due to the hypothesis $Q_{k}^{{A_{L}}}\leq Q_{k}^{A}$ and the fact that the matrix $I+\alpha\gamma DP\Pi_{Q{*}}-\alpha D$ is nonnegative, i.e., all elements are nonnegative by Lemma 4. The proof of the second inequality follows lines similar to the first proof. This completes the proof.
@@ C.3 Convergence of the lower comparison system
The lower comparison system in (5.3) can be divided into the linear parts with stochastic noises, $(I+\alpha\gamma DP\Pi_{Q^{*}}-\alpha D)(Q_{k}^{A_{L}}-Q^{*})+\alpha w_{k}^{A}$ and $(I+\alpha\gamma DP\Pi_{Q^{*}}-\alpha D)(Q_{k}^{B_{L}}-Q^{*})+\alpha w_{k}^{B}$, and the external disturbance parts, $\alpha\gamma DP(\Pi_{Q_{k}^{B}}-\Pi_{Q^{*}})(Q_{k}^{A}-Q_{k}^{B})$ and $\alpha\gamma DP(\Pi_{Q_{k}^{*}}-\Pi_{Q_{k}^{A}})(Q_{k}^{A}-Q_{k}^{B})$. As proved in B.4, the external disturbances are bounded. Using this fact, one can prove the finite-time error bounds of the linear part with stochastic noise as $(I+\alpha\gamma DP\Pi_{Q^{*}}-\alpha D)(Q_{k}^{A_{L}}-Q^{*})+\alpha w_{k}^{A}$ and $(I+\alpha\gamma DP\Pi_{Q^{*}}-\alpha D)(Q_{k}^{B_{L}}-Q^{*})+\alpha w_{k}^{B}$.
Theorem A.1.
@@
$\mathbb{E}[\lVert Q_{k}^{A_{L}}-Q^{*}\rVert_{\infty}]$ $\leq\frac{16\gamma d_{\mathrm{max}}\lvert\mathcal{S}\times\mathcal{A}\rvert\alpha^{1/2}}{d_{\mathrm{min}}^{7/2}(1-\gamma)^{9/2}}+\frac{24\rho^{k-3}k^{3}\lvert\mathcal{S}\times\mathcal{A}\rvert^{3/2}}{(1-\gamma)}+\frac{4\alpha^{1/2}\lvert\mathcal{S}\times\mathcal{A}\rvert}{d_{\mathrm{min}}^{1/2}(1-\gamma)^{3/2}}.$ (31)
@@
$Q_{k}^{A_{L}}-Q^{*}$ $=(I+\alpha\gamma DP\Pi_{Q^{*}}-\alpha D)^{k}\,(Q_{0}^{A_{L}}-Q^{*})+$
$\quad\underbrace{\sum_{j=0}^{k-1}\alpha\,(I+\alpha\gamma DP\Pi_{Q^{*}}-\alpha D)^{(k-1)-j}\,w_{j}^{A}}_{=:(*)}+\underbrace{\begin{aligned} \alpha\gamma DP\sum_{j=0}^{k-1}(I+\alpha\gamma DP\Pi_{Q^{*}}-\alpha D)^{(k-1)-j}\\ \quad\quad\times\,(\Pi_{Q_{j}^{B}}-\Pi_{Q^{*}})\,(Q_{j}^{A}-Q_{j}^{B})\end{aligned},}_{=:(**)}$
where ()(*) reflects the effect of the stochastic noise $w_{j}^{A}$ and ()(**) corresponds to the effect of the disturbance $Q_{j}^{A}-Q_{j}^{B}$. Taking the $\infty$-norm on the right-hand side of the above equation leads to
$\lVert Q_{k}^{A_{L}}-Q^{*}\rVert_{\infty}=$ $\biggl\lVert(I+\alpha\gamma DP\Pi_{Q^{*}}-\alpha D)^{k}(Q_{0}^{A_{L}}-Q^{*})+\sum_{j=0}^{k-1}\alpha(I+\alpha\gamma DP\Pi_{Q^{*}}-\alpha D)^{(k-1)-j}w_{j}^{A}$
$\quad+\alpha\gamma DP\sum_{j=0}^{k-1}(I+\alpha\gamma DP\Pi_{Q^{*}}-\alpha D)^{(k-1)-j}(\Pi_{Q_{j}^{B}}-\Pi_{Q^{*}})(Q_{j}^{A}-Q_{j}^{B})\biggr\rVert_{\infty}$
$\leq$ $\biggl\lVert(I+\alpha\gamma DP\Pi_{Q^{*}}-\alpha D)^{k}(Q_{0}^{A_{L}}-Q^{*})+\sum_{j=0}^{k-1}\alpha(I+\alpha\gamma DP\Pi_{Q^{*}}-\alpha D)^{(k-1)-j}w_{j}^{A}\biggr\rVert_{\infty}$
$\quad+\biggl\lVert\alpha\gamma DP\sum_{j=0}^{k-1}(I+\alpha\gamma DP\Pi_{Q^{*}}-\alpha D)^{(k-1)-j}(\Pi_{Q_{j}^{B}}-\Pi_{Q^{*}})(Q_{j}^{A}-Q_{j}^{B})\biggr\rVert_{\infty}$
$\leq$ $\underbrace{\biggl\lVert(I+\alpha\gamma DP{\Pi_{{Q^{*}}}}-\alpha D)^{k}(Q_{0}^{{A_{L}}}-{Q^{*}})}_{:=(*)}\underbrace{+\sum\limits_{j=0}^{k-1}\alpha{{(I+\alpha\gamma DP{\Pi_{{Q^{*}}}}-\alpha D)}^{(k-1)-j}}w_{j}^{A}\biggr\rVert_{\infty}}_{:=(**)}$
@@
where ()(*) and ()(**) in the second inequality corresponds to the solution of (16) with $x_{k}=Q_{k}^{{A_{L}}}-{Q^{*}}$ and $w_{k}=w_{k}^{A}$, we can apply the bound given in  Corollary A.2. Moreover, applying Lemma 7, one gets
$\lVert Q_{k}^{A_{L}}-Q^{*}\rVert_{\infty}\leq$ $\frac{4\alpha^{1/2}\lvert\mathcal{S}\times\mathcal{A}\rvert}{d_{\text{min}}^{1/2}(1-\gamma)^{3/2}}+\lvert\mathcal{S}\times\mathcal{A}\rvert^{3/2}\frac{2}{1-\gamma}\rho^{k}+\alpha\gamma d_{\text{max}}\sum_{j=0}^{k-1}\rho^{(k-1)-j}\biggl\lVert(\Pi_{Q_{j}^{B}}-\Pi_{Q^{*}})\biggr\rVert_{\infty}\biggl\lVert(Q_{j}^{A}-Q_{j}^{B})\biggr\rVert_{\infty},$
@@
$\mathbb{E}[\lVert Q_{k}^{A_{L}}-Q^{*}\rVert_{\infty}]$ $\leq\frac{8\gamma^{2}d_{\text{max}}^{2}\lvert\mathcal{S}\times\mathcal{A}\rvert\alpha^{1/2}}{d_{\text{min}}^{7/2}(1-\gamma)^{9/2}}+\frac{8\gamma d_{\text{max}}\lvert\mathcal{S}\times\mathcal{A}\rvert\alpha^{1/2}}{d_{\text{min}}^{5/2}(1-\gamma)^{7/2}}$
$\quad+\rho^{k-1}\biggl(\frac{(k-1)k(2k-1)}{6}\biggl)\frac{4\rho^{-2}\alpha^{2}\gamma^{2}d_{\text{max}}^{2}\lvert\mathcal{S}\times\mathcal{A}\rvert^{3/2}}{(1-\gamma)}$
$\quad+\rho^{k-1}\biggl(\frac{(k-1)k}{2}\biggl)\frac{4\rho^{-1}\alpha\gamma d_{\text{max}}\lvert\mathcal{S}\times\mathcal{A}\rvert^{3/2}}{(1-\gamma)}+\frac{4\alpha^{1/2}\lvert\mathcal{S}\times\mathcal{A}\rvert}{d_{\text{min}}^{1/2}(1-\gamma)^{3/2}}+\frac{2\rho^{k}\lvert\mathcal{S}\times\mathcal{A}\rvert^{3/2}}{(1-\gamma)}.$ (32)
@@
$\frac{8\gamma^{2}d_{\text{max}}^{2}\lvert\mathcal{S}\times\mathcal{A}\rvert\alpha^{1/2}}{d_{\text{min}}^{7/2}(1-\gamma)^{9/2}}+\frac{8\gamma d_{\text{max}}\lvert\mathcal{S}\times\mathcal{A}\rvert\alpha^{1/2}}{d_{\text{min}}^{5/2}(1-\gamma)^{7/2}}\leq 2\biggl(\frac{8\gamma d_{\text{max}}\lvert\mathcal{S}\times\mathcal{A}\rvert\alpha^{1/2}}{d_{\text{min}}^{7/2}(1-\gamma)^{9/2}}\biggl)$